# An Improved Lower Bound on the Image of the 2-adic Character Map for the Heisenberg Algebra via Modular Linear Differential Equations

### v1

@@ 1. Introduction
The study of $p$-adic vertex operator algebras was initiated by Franc-Mason [FM23b] in part as a means to further incorporate the $p$-adic study of modular forms into VOA theory. In particular, the normalized character map (i.e. 1-point correlation function or graded trace) on these new structures was shown to have image inside Serre’s space of $p$-adic modular forms [Ser73]. In the algebraic rather than the $p$-adic case, it is known that the normalized character map $F_{S}$ on the Heisenberg VOA $S$ gives a surjection of graded linear spaces $S\to M^{\prime}$ where $M^{\prime}$ denotes the space of quasi-modular forms — see [MT10] for a proof. Whether or not the $p$-adic character map surjects onto the ring of $p$-adic modular forms is currently an open question.
For the $p$-adic completion of $S$, which we denote as $S_{1}$, it was shown in [BF24] that the image of the character map contains families of $p$-adic modular forms, including every $p$-adic Eisenstein series $G_{k}^{\star}(q)$. In fact, a similar result was proven for a $p$-adic completion of a lattice vertex operator algebra, though in this manuscript we focus only on $S_{1}$. Then in [FM23a], the authors used the description of the space of $2$-adic modular forms as the Tate ring $\mathcal{M}_{2}=\mathbb{Q}_{2}\langle j^{-1}\rangle$ in the variable $j^{-1}$, where $j$ is the classical $j$-invariant, to prove that the space $\mathcal{M}_{2}^{\dagger}(7/4)$ of overconvergent modular forms of weight zero lies in the image of the character map on $S_{1}$. See Section 2.3 for a brief introduction to $p$-adic and overconvergent modular forms, and in particular for a definition of the space $M^{\dagger}_{2}(r)$. In Remark 6.8 of [FM23a], it was proposed that instead of working with the description $\mathcal{M}_{2}=\mathbb{Q}_{2}\langle j^{-1}\rangle$, the alternative description $\mathcal{M}_{2}=\mathbb{Q}_{2}\langle\lambda\rangle$, where $\lambda$ is a genus-zero Hauptmodul on $\Gamma_{0}(2)$, might yield better $p$-adic properties. This would be in analogy with similar computations in number theory, cf. [BC05, BK05]. Indeed, our first main theorem is the following:
@@
Since there are containments $\mathcal{M}_{2}^{\dagger}(r)\subseteq\mathcal{M}_{2}^{\dagger}(s)\subseteq\mathcal{M}_{2}$ when $s\leq r$, Theorem 1.1 represents an improvement over the analogous result of [FM23a].
@@
Now in all of [FM22, BF24, FM23a], the character on $S_{1}$ is defined via $p$-adic completion, and it is in particular the linear map $\widehat{F}_{S}$ making the following diagram commute:
@@
With Theorem 1.3 in hand, and knowing that the corresponding differential equation is generalized hypergeometric, we solve the MLDE using methods found in [FM16a, KT01], and this solution allows us to write our modular forms in the Eisenstein basis $\lambda_{n,m}\in\mathbb{Z}[E_{4},E_{6}]$ for all $n,m\geq 1$. With such explicit descriptions in hand, we can then use the formula of Mason-Tuite to lift $\lambda_{n,m}$ to the Heisenberg algebra. We then conclude the proof of Theorem 1.1 by studying the $p$-adic properties of these explicit formulas for the pre-images.
@@ 2.2. Modular Differential Operators
(2.10) $\mathcal{M}_{2}^{\dagger}(r)=\Set{a_{0}+a_{1}\lambda+a_{2}\lambda^{2}+\cdots}{|a_{n}|_{2}\cdot 2^{12nr}\to 0}.$
@@
$\mathcal{M}_{2}^{\dagger}(1/2)=\Set{a_{0}+a_{1}\widehat{F}_{S}(v_{1},q)+a_{2}\widehat{F}_{S}(v_{2},q)+\cdots}{|a_{n}|_{2}\cdot 2^{6n}\to 0}$
@@ 2.4. The Heisenberg VOA
$F_{S}(v,q)=\eta(q)Z_{S}(v,q)=\overline{\eta}(q)\sum_{n\geq 0}\text{Tr}_{S_{(n)}}o(v)q^{n}$
where $\overline{\eta}(q)=\prod_{i\geq 1}(1-q^{n})$ and $o(v)$ is the unique endomorphism associated to $v$ which preserves the round-bracket grading of $S$. We use the following re-formulation of equation (44) of [MT10] in order to compute $F_{S}(v,q)$ for a state $v=h[-k_{1}]h[-k_{2}]\cdots h[-k_{r}]\textbf{1}$ for $k_{i}\geq 1$ not necessarily distinct:
@@ 2.5. The p-adic Heisenberg VOA
We briefly recall some notions on $p$-adic VOAs also stated in [BF24], however the inclined reader should consult [FM23b] for all of the necessary details. In general, given an algebraic VOA $U$ over $\mathbb{Z}_{p}$, one can define the $\mathbb{Z}_{p}$-module
@@ 3.1. Serre’s Sequence
Let $p\in\{2,3,5,7,13\}$. Following Théorème 10 of [Ser73] and for any $n\geq 1$, we define a sequence $\lambda_{n,m}$ satisfying $\lambda^{n}=\lim_{m\to\infty}\lambda_{n,m}$. Fix $a,k\in\mathbb{N}$ such that $a\geq 1$ and $k\geq 4$ is even. Define
(3.1) $\lambda_{n,m}=\lambda^{n}g^{a\cdot p^{m}}+p^{1-\frac{ak}{2}\cdot p^{m}}\left.\left(\left.\lambda^{n}g^{a\cdot p^{m}}\right\rvert_{ak\cdot p^{m}}\hskip 2.0ptW\right)\hskip 2.0pt\right\rvert\hskip 1.0ptU_{p}.$
Note that $\lambda_{n,m}$ has weight $ak\cdot p^{m}$, the form $g\in M_{k}(\Gamma_{0}(p))$ is given by
$g=E_{k}(q)-p^{k}\cdot E_{k}(q^{p}),$
@@
$\left.\lambda^{n}g^{a\cdot p^{m}}\right\rvert_{ak\cdot p^{m}}\hskip 2.0ptW=p^{\frac{ak}{2}\cdot p^{m}}\left(p\tau\right)^{-ak\cdot p^{m}}\left(\frac{\eta(-1/\tau)}{\eta(-1/p\tau)}\right)^{\frac{24n}{p-1}}\left(E_{k}(-1/p\tau)-p^{k}E_{k}(-1/\tau)\right)^{a\cdot p^{m}},$
@@
$\left.\lambda^{n}g^{a\cdot p^{m}}\right\rvert_{ak\cdot p^{m}}\hskip 2.0ptW=p^{\frac{ak}{2}\cdot p^{m}-\frac{12n}{p-1}}\cdot\lambda^{-n}\cdot\left(E_{k}(p\tau)-E_{k}(\tau)\right)^{a\cdot p^{m}}$
@@
(3.2) $\lambda_{n,m}=\lambda^{n}\left(E_{k}(q)-p^{k}E_{k}(q^{p})\right)^{a\cdot p^{m}}+p^{1-\frac{12n}{p-1}}\left.\left(\lambda^{-n}\left(E_{k}(q^{p})-E_{k}(q)\right)^{a\cdot p^{m}}\right)\right\rvert\hskip 1.0ptU_{p}.$
@@
Remark 3.2.
From above, the sequence terms $\lambda_{n,m}$ seem to possess a pole of order $n$ at the zero cusp, which we clear by multiplying by $\Delta(q)^{n}$. That is, we consider the sequence $\Delta(q)^{n}\lambda_{n,m}\in M_{ak\cdot 2^{m}+12n}$ satisfying that $\Delta(q)^{n}\lambda^{n}=\lim_{m\to\infty}\Delta(q)^{n}\lambda_{n,m}$. Later, we will see that the $\lambda_{n,m}$ are modular of level one, and we will divide by $\Delta(q)^{n}$.
@@ 3.2. Sequences of MLDEs
$\lambda^{-n}\mathcal{E}_{4}(q)^{a\cdot 2^{m}}=\eta(q^{2})^{16a\cdot 2^{m}-24n}\eta(q)^{24n-8a\cdot 2^{m}}.$
@@
$\left.\left(\lambda^{-n}\mathcal{E}_{4}(q)^{3\cdot 2^{m}}\right)\right\rvert\hskip 1.0ptU_{2}=\eta(q)^{16a\cdot 2^{m}-24n}\left.\left(\eta(q)^{24n-8a\cdot 2^{m}}\right)\right\rvert\hskip 1.0ptU_{2}.$
Putting everything together and using identities derived in Lemma 3.3, eq. (3.2) with $k=4$ becomes
(3.5) $\lambda_{n,m}=\lambda^{n}\left(5E_{4}(q)-20E_{2}^{\star}(q)^{2}\right)^{a\cdot 2^{m}}+2^{1-12n}(240)^{a\cdot 2^{m}}\eta(q)^{16a\cdot 2^{m}-24n}\left.\left(\eta(q)^{24n-8a\cdot 2^{m}}\right)\right\rvert\hskip 1.0ptU_{2}.$
@@
It may be possible to write (3.5) explicitly without the $U_{2}$-operator, using the recursive formula given in page 5 of [BC05]. Since it is interesting enough that the $\Delta(q)^{n}\lambda_{n,m}$ will satisfy a “nice” MLDE, we keep the expression for the $\lambda_{n,m}$ as it is written above.
@@
Set $p=2$. For $n,m\geq 1$, the modular form $\Delta(q)^{n}\lambda_{n,m}$ with $k=4$ and $a\geq 1$ satisfies the third-degree MLDE:
(3.6) $\mathcal{D}^{3}(\Delta(q)^{n}\lambda_{n,m})=\left(\frac{3}{4}t_{n,m,a}^{2}+\frac{1}{4}t_{n,m,a}+\frac{1}{18}\right)\mathcal{D}(\Delta(q)^{n}\lambda_{n,m})E_{4}(q)\\ -\frac{1}{4}\left(t_{n,m,a}^{3}+t_{n,m,a}^{2}\right)\Delta(q)^{n}\lambda_{n,m}E_{6}(q),$
where $t_{n,m,a}=(1/3)(a\cdot 2^{m})-n$.
@@
$\lambda_{n,m}$ $=\lambda^{n}X^{a\cdot 2^{m}}+2^{1-12n}(240)^{a\cdot 2^{m}}\eta(q)^{16a\cdot 2^{m}-24n}\left.\left(\eta(q)^{24n-8a\cdot 2^{m}}\right)\right\rvert\hskip 1.0ptU_{2}$
$:=T_{1}+2^{1-12n}(240)^{a\cdot 2^{m}}T_{2}.$
@@
(3.7) $\mathcal{D}(T_{1})=t_{n,m,a}\lambda^{n}E_{2}^{\star}(q)X^{a\cdot 2^{m}}$
@@
$\mathcal{D}^{2}(T_{1})=\lambda^{n}X^{a\cdot 2^{m}}\left(\left(t_{n,m,a}^{2}+\frac{1}{3}t_{n,m,a}\right)E_{2}^{\star}(q)^{2}-\frac{1}{6}_{n,m,a}E_{4}(q)\right).$
@@
$\mathcal{D}^{3}(T_{1})=\lambda^{n}X^{a\cdot 2^{m}}\left(\left(t_{n,m,a}^{3}+t_{n,m,a}^{2}\right)E_{2}^{\star}(q)^{3}+\left(\frac{1}{18}t_{n,m,a}-\frac{1}{2}t_{n,m,a}^{2}\right)E_{2}^{\star}(q)E_{4}(q)\right).$
@@
$\mathcal{D}^{3}(T_{1})=\lambda^{n}X^{a\cdot 2^{m}}\left(\left(\frac{3}{4}t_{n,m,a}^{3}+\frac{1}{4}t_{n,m,a}^{2}+\frac{1}{18}t_{n,m,a}\right)E_{2}^{\star}(q)E_{4}(q)\right.\\ \left.-\frac{1}{4}\left(t_{n,m,a}^{3}+t_{n,m,a}^{2}\right)\left(-4E_{2}^{\star}(q)^{3}+3E_{2}^{\star}(q)E_{4}(q)\right)\right)$
@@
$\mathcal{D}^{i}(T_{2})=\eta(q)^{16a\cdot 2^{m}-24n}\mathcal{D}^{i}\left(\left.\eta(q)^{-24c_{n,m,a}}\right\rvert\hskip 1.0ptU_{2}\right)$
@@
$\mathcal{D}^{2}\left(\left.\eta(q)^{-24t_{n,m,a}}\right\rvert\hskip 1.0ptU_{2}\right)=\left.\left(\eta(q)^{-24t_{n,m,a}}\left(\left(\frac{t_{n,m,a}^{2}}{4}-\frac{t_{n,m,a}}{8}\right)E_{2}^{\star}(q)^{2}+\frac{t_{n,m,a}}{24}E_{4}(q)\right)\right)\right\rvert\hskip 1.0ptU_{2},$
@@
$\mathcal{D}^{3}\left(\left.\eta(q)^{-24t_{n,m,a}}\right\rvert\hskip 1.0ptU_{2}\right)=\left.\left(\eta(q)^{-24t_{n,m,a}}\left(\left(-\frac{t_{n,m,a}^{3}}{8}+\frac{3t_{n,m,a}^{2}}{16}-\frac{5t_{n,m,a}}{144}\right)E_{2}^{\star}(q)^{3}\right.\right.\right.\\ +\left.\left.\left.\left(-\frac{t_{n,m,a}^{2}}{16}+\frac{t_{n,m,a}}{144}\right)E_{2}^{\star}(q)E_{4}(q)\right)\right)\right\rvert\hskip 1.0ptU_{2}.$
@@
$\mathcal{D}^{3}\left(\left.\eta(q)^{-24t_{n,m,a}}\right\rvert\hskip 1.0ptU_{2}\right)=\left.\left(\left(-\frac{15}{32}t_{n,m,a}^{3}-\frac{5}{32}t_{n,m,a}^{2}-\frac{5}{144}t_{n,m,a}\right)\eta(q)^{-24t_{n,m,a}}E_{2}^{\star}(q)^{3}\right.\right.\\ \left.\left.+\left(\frac{3}{32}t_{n,m,a}^{3}+\frac{1}{32}t_{n,m,a}^{2}+\frac{1}{144}t_{n,m,a}\right)\eta(q)^{-24t_{n,m,a}}E_{2}^{\star}(q)E_{4}(q)\right.\right.\\ \left.\left.-\frac{3}{32}\left(t_{n,m,a}^{3}+t_{n,m,a}^{2}\right)\eta(q)^{-24t_{n,m,a}}E_{2}^{\star}(q)E_{4}(q)\right.\right.\\ \left.\left.+\frac{11}{32}\left(t_{n,m,a}^{3}+t_{n,m,a}^{2}\right)\eta(q)^{-24t_{n,m,a}}E_{2}^{\star}(q)^{3}\right)\right\rvert\hskip 1.0ptU_{2}$
@@
$\mathcal{D}^{3}\left(\left.\eta(q)^{-24t_{n,m,a}}\right\rvert\hskip 1.0ptU_{2}\right)=\left(\frac{3}{4}t_{n,m,a}^{2}+\frac{1}{4}t_{n,m,a}+\frac{1}{18}\right)\mathcal{D}\left(\left.\eta(q)^{-24t_{n,m,a}}\right\rvert\hskip 1.0ptU_{2}\right)E_{4}(q)\\ -\frac{1}{4}\left(t_{n,m,a}^{3}+t_{n,m,a}^{2}\right)\left(\left.\eta(q)^{-24t_{n,m,a}}\right\rvert\hskip 1.0ptU_{2}\right)E_{6}(q)$
@@
The MLDE given in the statement of Lemma 3.3 is the $2$-adic limit of the sequence of MLDEs whose $m$-th term is given in Proposition 3.5. This can easily be shown by writing eq. (3.3) in the basis $\mathbb{C}[E_{2}^{\star},E_{4}]$ of $M(\Gamma_{0}(2))$, using the expression for $\mathcal{D}(\lambda^{n})$ and noting that $\lim_{m\to\infty}c_{n,m,a}^{r}=(-n)^{r}$.
@@
$\lambda_{n,m}=\frac{\eta(q^{3})^{12n}}{\eta(q)^{12n}}\cdot(-80)^{a\cdot 3^{m}}\left(\left(\frac{\eta(q)^{3}}{\eta(q^{3})}\right)^{4}+9q\cdot\left(\frac{\eta(q)^{9}}{\eta(q)}\right)^{3}\left(\frac{\eta(q)^{3}}{\eta(q^{3})}\right)^{4}\right)^{a\cdot 3^{m}}\\ +3^{1-6n}\cdot(-240)^{a\cdot 3^{m}}\left(\frac{\eta(q)^{12n}}{\eta(q^{3})^{12n}}\right)\left.\left(\eta(q^{3})^{8}+9\eta(q^{3})^{8}\left(\frac{\eta(q^{9})}{\eta(q)}\right)^{3}\right)^{a\cdot 3^{m}}\right\rvert\hskip 1.0ptU_{3}$

### v2

@@ 1. Introduction
The study of $p$-adic vertex operator algebras was initiated by Franc-Mason [FM23] in part as a means to further incorporate the $p$-adic study of modular forms into VOA theory. In particular, the normalized character map (i.e. 1-point correlation function or graded trace) on these new structures was shown to have image inside Serre’s space of $p$-adic modular forms [Ser73]. In the algebraic rather than the $p$-adic case, it is known that the normalized character map $F_{S}$ on the Heisenberg VOA $S$ gives a surjection of graded linear spaces $S\to M^{\prime}$ where $M^{\prime}$ denotes the space of quasi-modular forms — see [MT10] for a proof. Whether or not the $p$-adic character map surjects onto the ring of $p$-adic modular forms is currently an open question.
For the $p$-adic completion of $S$, which we denote as $S_{1}$, it was shown in [BF24] that the image of the character map contains families of $p$-adic modular forms, including every $p$-adic Eisenstein series $G_{k}^{\star}(q)$. In fact, a similar result was proven for a $p$-adic completion of a lattice vertex operator algebra, though in this manuscript we focus only on $S_{1}$. Then in [FM25], the authors used the description of the space of $2$-adic modular forms as the Tate ring $\mathcal{M}_{2}=\mathbb{Q}_{2}\langle j^{-1}\rangle$ in the variable $j^{-1}$, where $j$ is the classical $j$-invariant, to prove that the space $\mathcal{M}_{2}^{\dagger}(7/4)$ of overconvergent modular forms of weight zero lies in the image of the character map on $S_{1}$. See Section 2.3 for a brief introduction to $p$-adic and overconvergent modular forms, and in particular for a definition of the space $M^{\dagger}_{2}(r)$. In Remark 6.8 of [FM25], it was proposed that instead of working with the description $\mathcal{M}_{2}=\mathbb{Q}_{2}\langle j^{-1}\rangle$, the alternative description $\mathcal{M}_{2}=\mathbb{Q}_{2}\langle\lambda\rangle$, where $\lambda$ is a genus-zero Hauptmodul on $\Gamma_{0}(2)$, might yield better $p$-adic properties. This would be in analogy with similar computations in number theory, cf. [BC05, BK05]. Indeed, our first main theorem is the following:
@@
Since there are containments $\mathcal{M}_{2}^{\dagger}(r)\subseteq\mathcal{M}_{2}^{\dagger}(s)\subseteq\mathcal{M}_{2}$ when $s\leq r$, Theorem 1.1 represents an improvement over the analogous result of [FM25].
@@
Now in all of [FM22, BF24, FM25], the character on $S_{1}$ is defined via $p$-adic completion, and it is in particular the linear map $\widehat{F}_{S}$ making the following diagram commute:
@@
With Theorem 1.3 in hand, and knowing that the corresponding differential equation is generalized hypergeometric, we solve the MLDE using methods found in [FM16, KT01], and this solution allows us to write our modular forms in the Eisenstein basis $\lambda_{n,m}\in\mathbb{Z}[E_{4},E_{6}]$ for all $n,m\geq 1$. With such explicit descriptions in hand, we can then use the formula of Mason-Tuite to lift $\lambda_{n,m}$ to the Heisenberg algebra. We then conclude the proof of Theorem 1.1 by studying the $p$-adic properties of these explicit formulas for the pre-images.
@@
The authors would also like to thank the referee for their helpful comments on a prior version of this manuscript.
@@ 2.2. Modular Differential Operators
(2.10) $\mathcal{M}_{2}^{\dagger}(r)=\mathbb{Q}_{2}\langle 2^{12r}\lambda\rangle=\Set{a_{0}+a_{1}\lambda+a_{2}\lambda^{2}+\cdots}{|a_{n}|_{2}\cdot 2^{12nr}\to 0}.$
@@
$\mathcal{M}_{2}^{\dagger}(1/2)=\mathbb{Q}_{2}\langle 2^{6}\lambda\rangle=\Set{a_{0}+a_{1}\widehat{F}_{S}(v_{1},q)+a_{2}\widehat{F}_{S}(v_{2},q)+\cdots}{|a_{n}|_{2}\cdot 2^{6n}\to 0}$
@@ 2.4. The Heisenberg VOA
$F_{S}(v,q)=\eta(q)Z_{S}(v,q)=q^{-1/24}\eta(q)\sum_{n\geq 0}\text{Tr}_{S_{(n)}}o(v)q^{n}$
where $o(v)$ is the unique endomorphism associated to $v$ which preserves the round-bracket grading of $S$. Note above that the trace is taken over the round-bracket grading. We use the following re-formulation of equation (44) of [MT10] in order to compute $F_{S}(v,q)$ for a state $v=h[-k_{1}]h[-k_{2}]\cdots h[-k_{r}]\textbf{1}$ for $k_{i}\geq 1$ not necessarily distinct:
@@ 2.5. The p-adic Heisenberg VOA
We briefly recall some notions on $p$-adic VOAs also stated in [BF24], however the inclined reader should consult [FM23] for all of the necessary details. In general, given an algebraic VOA $U$ over $\mathbb{Z}_{p}$, one can define the $\mathbb{Z}_{p}$-module
@@ 3.1. Serre’s Sequence
Let $p\in\{2,3,5,7,13\}$. Following Théorème 10 of [Ser73] and for any $n\geq 1$, we construct a sequence $\lambda_{n,m}$ of modular forms of level one satisfying $\lambda^{n}=\lim_{m\to\infty}\lambda_{n,m}$ by tracing over the representatives of $\Gamma_{0}(p)/\text{SL}_{2}(\mathbb{Z})$. We will see shortly that the $\lambda_{n,m}$ possess a pole of order $n$ at the zero cusp, hence we will be considering the sequence $\Delta(q)^{n}\lambda_{n,m}\in M_{12\cdot 2^{m}+12n}$ satisfying that $\Delta(q)^{n}\lambda^{n}=\lim_{m\to\infty}\Delta(q)^{n}\lambda_{n,m}$. Only after realizing each as the solution of an MLDE in Proposition 3.4 and getting explicit expressions in Proposition 3.7 will we see that the $\lambda_{n,m}$ are modular of level one as expected, and we will divide by $\Delta(q)^{n}$. Let us set
[…]
Note that $\lambda_{n,m}$ has weight $12\cdot p^{m}$, the form $g\in M_{4}(\Gamma_{0}(p))$ is given by
$g=E_{4}(q)-p^{4}\cdot E_{4}(q^{p}),$
@@
$\left.\lambda^{n}g^{3\cdot p^{m}}\right\rvert_{12\cdot p^{m}}\hskip 2.0ptW=p^{6\cdot p^{m}}\left(p\tau\right)^{-12\cdot p^{m}}\left(\frac{\eta(-1/\tau)}{\eta(-1/p\tau)}\right)^{\frac{24n}{p-1}}\left(E_{4}(-1/p\tau)-p^{4}E_{4}(-1/\tau)\right)^{3\cdot p^{m}},$
@@
$\left.\lambda^{n}g^{3\cdot p^{m}}\right\rvert_{12\cdot p^{m}}\hskip 2.0ptW=p^{6\cdot p^{m}-\frac{12n}{p-1}}\cdot\lambda^{-n}\cdot\left(E_{4}(p\tau)-E_{4}(\tau)\right)^{3\cdot p^{m}}$
@@
(3.2) $\lambda_{n,m}=\lambda^{n}\left(E_{4}(q)-p^{4}E_{4}(q^{p})\right)^{3\cdot p^{m}}+p^{1-\frac{12n}{p-1}}\left.\left(\lambda^{-n}\left(E_{4}(q^{p})-E_{4}(q)\right)^{3\cdot p^{m}}\right)\right\rvert\hskip 1.0ptU_{p}.$
@@ 3.2. Sequences of MLDEs
In this subsection, we formulate the explicit MLDEs needed to write the $\lambda_{n,m}$ in terms of an Eisenstein basis. In fact within Proposition 3.4, we prove Theorem 1.3.
@@
$\lambda^{-n}\mathcal{E}_{4}(q)^{3\cdot 2^{m}}=\Delta(q^{2})^{2^{m+1}-n}\Delta(q)^{n-2^{m}}.$
@@
$\left.\left(\lambda^{-n}\mathcal{E}_{4}(q)^{3\cdot 2^{m}}\right)\right\rvert\hskip 1.0ptU_{2}=\Delta(q)^{2^{m+1}-n}\left.\left(\Delta(q)^{n-2^{m}}\right)\right\rvert\hskip 1.0ptU_{2}.$
Putting everything together and using identities derived in Lemma 3.2, eq. (3.2) becomes
(3.5) $\lambda_{n,m}=\lambda^{n}\left(5E_{4}(q)-20E_{2}^{\star}(q)^{2}\right)^{3\cdot 2^{m}}+2^{1-12n}(240)^{3\cdot 2^{m}}\Delta(q)^{2^{m+1}-n}\left.\left(\Delta(q)^{n-2^{m}}\right)\right\rvert\hskip 1.0ptU_{2}.$
@@
It may be possible to write (3.5) explicitly without the $U_{2}$-operator, using the recursive formula given in page 5 of [BC05] or methods in [Des14]. Since it is interesting enough that the $\Delta(q)^{n}\lambda_{n,m}$ will satisfy a “nice” MLDE, we keep the expression for the $\lambda_{n,m}$ as it is written above.
@@
Set $p=2$. For $n,m\geq 1$, the modular form $\Delta(q)^{n}\lambda_{n,m}$ satisfies the third-degree MLDE:
(3.6) $\mathcal{D}^{3}(\Delta(q)^{n}\lambda_{n,m})=\left(\frac{3}{4}t_{n,m}^{2}+\frac{1}{4}t_{n,m}+\frac{1}{18}\right)\mathcal{D}(\Delta(q)^{n}\lambda_{n,m})E_{4}(q)\\ -\frac{1}{4}\left(t_{n,m}^{3}+t_{n,m}^{2}\right)\Delta(q)^{n}\lambda_{n,m}E_{6}(q),$
where $t_{n,m}=2^{m}-n$.
@@
$\lambda_{n,m}$ $=\lambda^{n}X^{3\cdot 2^{m}}+2^{1-12n}(240)^{3\cdot 2^{m}}\Delta(q)^{2^{m+1}-n}\left.\left(\Delta(q)^{n-2^{m}}\right)\right\rvert\hskip 1.0ptU_{2}$
$:=T_{1}+2^{1-12n}(240)^{3\cdot 2^{m}}T_{2}.$
@@
(3.7) $\mathcal{D}(T_{1})=t_{n,m}\lambda^{n}E_{2}^{\star}(q)X^{3\cdot 2^{m}}$
@@
$\mathcal{D}^{2}(T_{1})=\lambda^{n}X^{3\cdot 2^{m}}\left(\left(t_{n,m}^{2}+\frac{1}{3}t_{n,m}\right)E_{2}^{\star}(q)^{2}-\frac{1}{6}t_{n,m}E_{4}(q)\right).$
@@
$\mathcal{D}^{3}(T_{1})=\lambda^{n}X^{3\cdot 2^{m}}\left(\left(t_{n,m}^{3}+t_{n,m}^{2}\right)E_{2}^{\star}(q)^{3}+\left(\frac{1}{18}t_{n,m}-\frac{1}{2}t_{n,m}^{2}\right)E_{2}^{\star}(q)E_{4}(q)\right).$
@@
$\mathcal{D}^{3}(T_{1})=\lambda^{n}X^{3\cdot 2^{m}}\left(\left(\frac{3}{4}t_{n,m}^{3}+\frac{1}{4}t_{n,m}^{2}+\frac{1}{18}t_{n,m}\right)E_{2}^{\star}(q)E_{4}(q)\right.\\ \left.-\frac{1}{4}\left(t_{n,m}^{3}+t_{n,m}^{2}\right)\left(-4E_{2}^{\star}(q)^{3}+3E_{2}^{\star}(q)E_{4}(q)\right)\right)$
@@
$\mathcal{D}^{i}(T_{2})=\Delta(q)^{2^{m+1}-n}\mathcal{D}^{i}\left(\left.\Delta(q)^{-t_{n,m}}\right\rvert\hskip 1.0ptU_{2}\right)$
@@
[…]
@@
$\mathcal{D}^{3}\left(\left.\Delta(q)^{-t_{n,m}}\right\rvert\hskip 1.0ptU_{2}\right)=\left.\left(\Delta(q)^{-t_{n,m}}\left(\left(-\frac{t_{n,m}^{3}}{8}+\frac{3t_{n,m}^{2}}{16}-\frac{5t_{n,m}}{144}\right)E_{2}^{\star}(q)^{3}\right.\right.\right.\\ +\left.\left.\left.\left(-\frac{t_{n,m}^{2}}{16}+\frac{t_{n,m}}{144}\right)E_{2}^{\star}(q)E_{4}(q)\right)\right)\right\rvert\hskip 1.0ptU_{2}.$
@@
$\mathcal{D}^{3}\left(\left.\Delta(q)^{-t_{n,m}}\right\rvert\hskip 1.0ptU_{2}\right)=\left.\left(\left(-\frac{15}{32}t_{n,m}^{3}-\frac{5}{32}t_{n,m}^{2}-\frac{5}{144}t_{n,m}\right)\Delta(q)^{-t_{n,m}}E_{2}^{\star}(q)^{3}\right.\right.\\ \left.\left.+\left(\frac{3}{32}t_{n,m}^{3}+\frac{1}{32}t_{n,m}^{2}+\frac{1}{144}t_{n,m}\right)\Delta(q)^{-t_{n,m}}E_{2}^{\star}(q)E_{4}(q)\right.\right.\\ \left.\left.-\frac{3}{32}\left(t_{n,m}^{3}+t_{n,m}^{2}\right)\Delta(q)^{-t_{n,m}}E_{2}^{\star}(q)E_{4}(q)\right.\right.\\ \left.\left.+\frac{11}{32}\left(t_{n,m}^{3}+t_{n,m}^{2}\right)\Delta(q)^{-t_{n,m}}E_{2}^{\star}(q)^{3}\right)\right\rvert\hskip 1.0ptU_{2}$
@@
$\mathcal{D}^{3}\left(\left.\Delta(q)^{-t_{n,m}}\right\rvert\hskip 1.0ptU_{2}\right)=\left(\frac{3}{4}t_{n,m}^{2}+\frac{1}{4}t_{n,m}+\frac{1}{18}\right)\mathcal{D}\left(\left.\Delta(q)^{-t_{n,m}}\right\rvert\hskip 1.0ptU_{2}\right)E_{4}(q)\\ -\frac{1}{4}\left(t_{n,m}^{3}+t_{n,m}^{2}\right)\left(\left.\Delta(q)^{-t_{n,m}}\right\rvert\hskip 1.0ptU_{2}\right)E_{6}(q)$
@@
The MLDE given in the statement of Lemma 3.2 is the $2$-adic limit of the sequence of MLDEs whose $m$-th term is given in Proposition 3.4. This can easily be shown by writing eq. (3.3) in the basis $\mathbb{C}[E_{2}^{\star},E_{4}]$ of $M(\Gamma_{0}(2))$, using the expression for $\mathcal{D}(\lambda^{n})$ and noting that $\lim_{m\to\infty}t_{n,m}^{r}=(-n)^{r}$.
@@
$\lambda_{n,m}=\frac{\eta(q^{3})^{12n}}{\eta(q)^{12n}}\cdot(-80)^{3^{m+1}}\left(\left(\frac{\eta(q)^{3}}{\eta(q^{3})}\right)^{4}+9q\cdot\left(\frac{\eta(q)^{9}}{\eta(q)}\right)^{3}\left(\frac{\eta(q)^{3}}{\eta(q^{3})}\right)^{4}\right)^{3^{m+1}}\\ +3^{1-6n}\cdot(-240)^{3^{m+1}}\left(\frac{\eta(q)^{12n}}{\eta(q^{3})^{12n}}\right)\left.\left(\eta(q^{3})^{8}+9\eta(q^{3})^{8}\left(\frac{\eta(q^{9})}{\eta(q)}\right)^{3}\right)^{3^{m+1}}\right\rvert\hskip 1.0ptU_{3}$